### 26 Feb 2024 v1

@@ 6 Multilayered patters
For all $n\in\mathbb{Z}^{+}$, for some $m\in\mathbb{Z}^{+}$ there exists a 7-colored $m\times m$ pattern, $p_{n}$, such that no aTAM system $\mathcal{T}_{\leq n}=(T,\sigma,\tau)$ where $|T|\leq n$, $|\sigma|=1$, and $\tau\leq 2$ weakly self-assembles $p_{n}$, but a barely-3DaTAM system $\mathcal{T}_{p_{n}}=(T_{p_{n}},\sigma_{p_{n}},2)$ where $|T_{p_{n}}|=O(\log{n}/\log{\log{n}})$ and $|\sigma_{p_{n}}|=1$ weakly self-assembles $p_{n}$.
Proof
[…]
@@
Figure 7: Example $p_{n}$ patterns created by the construction in the proof of Theorem 6.1. The (repeatedly copied) binary sequence of results from the simulations of aTAM systems starts at the top, with the two colors of that row, and all subsequent boundary rows, being determined by the first bit of that pattern. During the downward growth from that row, during which a full square is formed, the repeating grid formed by the copies of that pattern is copied both downward and to both sides. The boundary columns also have two colors determined by the first bit of the sequence (one of them the same as in the boundary rows) for a total of 3 boundary colors. The interiors always use the same 4 colors.
The bit sequence that determines the colors used by the boundaries, and the ordering of the colors on the boundaries and in the interiors, is determined via simulations of a series of aTAM systems. Intuitively, our proof utilizes a construction that performs a diagonalization against all possible aTAM systems with $\leq n$ tile types by simulating each, and for each keeping track of the color of tile it places in a location specific to the index of that system so that it can ultimately generate the colored pattern $p_{n}$ that differs in at least one location from every simulated system. The dimensions of each $c\times c$ cell are $c=T(n)$, where $T(n)$ is a function that takes a number of tile types and returns a count of all possible aTAM systems with $\leq n$ tile types (with the caveat that it “overcounts” by allowing for functionally duplicate systems, as this does not impact the correctness and makes the discussion of the proof simpler). The colors of the rows and columns encode the bit sequence generated by the simulations, with the same bit sequence encoded in both the rows and the columns via an assignment of colors. There is a unique color assigned for each intersection of two bits (i.e. 00, 01, 10, and 11), with 4 colors reserved for boundaries of grid cells and 4 separate colors reserved for the interior locations of the grid cells.
Figure 8: Schematic overview of the portion of the construction for the proof of Theorem 6.1 that grows in plane $z=0$. Note that the modules are not shown to scale. (Green) The seed tile, (Fuchsia) the base conversion module that unpacks the binary representation of $n$, (Aqua) the module that simulates the Turing machine $M$ on input $n$, that in-turn simulates each TAS with $\leq n$ tile types in sequence and saves a result bit (Black or White) for each, (Yellow and Grey) the pattern of result bits is copied repetitively to the right until it covers the entire top row. (Light Grey) A simple “filler” tile causes the assembly to form a complete square.
At a high-level, the construction consists of a handful of components (of varying complexity) that can be seen schematically depicted in Figure 8. The number $n$ is encoded in $O(\log{n}/\log{\log{n}})$ tile types following the technique of [1]. From the seed tile of $\mathcal{T}_{p_{n}}$, the $O(\log{n}/\log{\log{n}})$ tiles representing $n$ in an optimally compressed base grow to the right. Then, rows grow upward and to the right to do a base conversion in which the bits of $n$ are “unpacked” so that the northern glues of the tiles of top row of that triangle represent $n$ as $\log{n}$ bits (shown in fuchsia in Figure 8). (Figure 9 depicts a slightly more detailed example of the bit unpacking.) A simple set of filler tiles grow to the south of the seed’s row to form the bottom of the triangle. The tile complexity of this stage of the construction is $O(\log{n}/\log{\log{n}})$. The tile complexity of the remaining portions of the construction is $O(1)$, as they use a constant number of tile types independent of $n$.
Figure 9: Schematic example of the growth of the base conversion module’s growth from Figure 8. The seed tile is shown in green. Using the technique of [1], the number $n$ is encoded using $O(\log{n}/\log{\log{n}})$ tile types. These tile types form the (white) row that grows to the east of the seed. A base conversion then occurs via rows that grow to the north (fuchsia) to convert $n$ to binary, so that the northern row of this module (white) consists of tiles that encode $n$ in binary.
6.1 Simulation of all aTAM systems with $\leq n$ tile types
A zig-zag Turing machine module (i.e., a standard aTAM construction in which a growing assembly simulates a Turing machine while rows grow in alternating, zig-zag, directions and each row increases in length by one tile - see [10, 7, 2] for some examples) uses $n$ as input. The Turing machine $M$ being simulated runs the procedure SimulateAllTileAssemblySystems, with argument $n$, as shown in Algorithm 1 (and depicted in aqua in Figure 8).
SimulateAllTileAssemblySystems is given the number $n$ as the maximum number of tile types allowed in a system, and first uses that to compute the number of all possible aTAM tile sets that could be created with up to $n$ tile types. [^1] Note that it is only necessary to consider a single tile set of each set of equivalent tile sets. Two tile sets are considered equivalent if they are isomorphic to each other, meaning that there is a bijection between all glue and tile types, since equivalent tile sets would form the same pattern. The number of possible glues for each side of a tile (recognizing that having more glue labels than tile types is equivalent to having glues with no binding partners, which is equivalent to making those unmatched glues null) is $O(n)$, and with 8 possible colors and 4 sides for any tile type, the number of possible tile types is $O(n^{4})$. To determine all possible aTAM systems that could be made from this set of possible tile types, we overcount (since it does not change the correctness and it makes the analysis and proof easier to follow) as $O((n^{4})^{n^{4}})=O(n^{4n^{4}})$. Note that this accounts for all tile sets of size $\leq n$ since one of the tile types is the null tile type with no glues (which can simply be omitted from any set without changing its behavior) while also counting many identical (i.e. functionally equivalent) tile sets multiple times. Additionally, while all of the simulations will be at temperature $\tau=2$, since all possible tile sets will be generated with only strength-2 glues, and when simulated at $\tau=2$ these are functionally equivalent to tile sets where the glues are changed to strength-1 and simulated at $\tau=1$, this method of counting causes the behaviors of all possible systems with $\leq n$ tile types and $\tau\leq 2$ to be simulated (again, many of them multiple times).
To generate all possible aTAM systems with single-tile seeds from that collection of possible tile sets, since we can keep the temperature at $\tau=2$ (as previously explained), we only need to vary the tile types that serve as the seed (because we only consider systems with single-tile seeds). Creating a unique system for each tile type in a tile set to be the seed yields $O(n*n^{4n^{4}})=O(n^{4n^{4}+1})$ aTAM systems. For simplicity, we refer to this value as $T(n)$ (as previously mentioned). Since each of the $T(n)$ systems will be simulated and a bit value will be saved for each (ultimately allowing the construction to build a pattern $p_{n}$ that differs from each of them), that bit sequence will be of length $T(n)$. We will use the value of $T(n)$ to determine the number of steps for which each system must be simulated.
Recall that $p_{n}$ consists of a grid of cells of size $T(n)\times T(n)$ repeated horizontally and vertically. As each system $\mathcal{T}_{i}$, for $0<i\leq T(n)$, is simulated, its index $i$ is noted so that the construction can guarantee that the $i$th row and $i$th column of each cell will differ from a (potentially) corresponding location in the assembly produced by $\mathcal{T}_{i}$ during its simulation. We want to ensure that for each $\mathcal{T}_{i}$, if it happens to make a repeating grid of $T(n)\times T(n)$ cells composed of 7 colors, $p_{n}$ differs in at least one location of each cell from at least one location of one cell of the pattern produced by $\mathcal{T}_{i}$. Any system $\mathcal{T}_{i}$ that does not even produce a single valid cell of size $T(n)\times T(n)$ bounded by the boundary colors has no chance of generating $p_{n}$ so can be easily discounted. For all other $T_{i}$, a bit computed after running the simulation of $T_{i}$ is used to ensure $p_{n}$ differs.
6.2 Making $p_{n}$ differ from each simulated system
For the simulation of each system, we do not impose a restriction upon the translation of the pattern that the system makes relative to our target pattern $p_{n}$. Therefore, we cannot assume the relative position of the seed tile of $\mathcal{T}_{i}$ with respect to any portion of $p_{n}$, and we simulate each $\mathcal{T}_{i}$ until it grows an assembly that has at least one dimension (width or height) that spans the distance of a full cell with boundaries on both sides. To do so, we simulate each $\mathcal{T}_{i}$ for $4T(n)^{2}$ steps because this is the number of tiles contained within a $2\times 2$ square of grid cells, ensuring that irrespective of the position of the seed tile with respect to the pattern formed, the full dimension of at least one grid cell and its boundaries in that dimension must be spanned. Figure 10 shows an example of grid cells and a bounding box of that size, demonstrating why such bounds suffice. (Note that any system that becomes terminal before reaching such a size clearly cannot weakly self-assembly $p_{n}$, since it is much larger, so we can record an arbitrary output bit for that simulation.) Since we are only concerned with systems that can create grid-like patterns with the same boundary and interior colors used by $p_{n}$ (as differing patterns created by other systems will immediately disagree with $p_{n}$), we can inspect the assembly $\alpha_{i}$ produced by the simulation of $\mathcal{T}_{i}$ as follows.
Figure 10: A portion of a pattern $p_{n}$ for the bit sequence $11010101$, showing a $4\times 4$ grid of cells and a bounding box (Red) enclosing a $2\times 2$ portion of the grid, containing $4T(n)^{2}$ tiles, where $T(n)$ is the width and height of a cell. Any assembly containing $4T(n)^{2}$ tiles must contain a connected component of at least width $\geq T(n)+1$ and/or height $\geq T(n)+1$, ensuring two boundary locations on the sides of an assembly spanning a cell.
@@ 6.2 Making $p_{n}$ differ from each simulated system
Figure 11: An example assembly $\alpha_{i}$ possibly formed during the simulation of some $T_{i}$. The leftmost red square hightlights a leftmost tile with a boundary color, and the corresponding black rectangle highlights the rest of that column and thus the boundary of a (potential) cell. Assuming $T(n)=8$, i.e., grid cell sizes of 8, the rightmost black rectangle highlights the locations of tiles at the boundary of the next cell to the right. Assuming an index value of $i=6$, the rightmost red square highlights a tile at that index, with respect to the boundary to the left. The color of the highlighted cell is Yellow, which is one of the two colors reserved for columns representing the bit value 0. Therefore, the bit value 1 is saved for index $i$ to ensure that the pattern $p_{n}$ will never place a Yellow tile $i$ locations to the east of a tile with a boundary color, guaranteeing that $p_{n}$ differs from the pattern produced by $T_{i}$.
The simulation of $M$ proceeds through the simulation of each of the $T(n)$ aTAM systems with $\leq n$ tile types. As the rows representing each simulation grow, they pass the currently computed sequence of bits upward through the tiles performing the simulations. Once $M$ completes, the sequence of generated bits is encoded in the north glues of the leftmost $T(n)$ tiles of the top row. This is represented in Figure 8 as the Black and White sequence on the top of the aqua-colored portion of the wedge. Note that the tile types that simulate $M$ are a constant-sized tile set, regardless of the value of $n$.
At that point, another constant-sized set of tiles grow in a zig-zag manner to copy the bit sequence over and over, to the right, until the entire top row consists of copies of the bit sequence (with the last copy of the sequence potentially truncated). Note that during all of the diagonal upward growth, a single “filler” tile type attaches to the right of the diagonal so that once the northward growth completes the full assembly will be a square.
Figure 12: An example portion of the assembly from the proof of Theorem 6.1 that shows the copying of the bit sequence $11010101$ to the right until it occupies the entire top row. This corresponds to the portion of the construction shown in Figure 8 above the aqua portion. At the completion of the bit sequence copying, the rightmost tile of the topmost row (shown here in red), initiates the growth into the second plane in which the pattern $p_{n}$ will assemble. The grey portion below the yellow is formed by a single “filler” tile that causes the final assembly to be a square.
[…]
Figure 13: The growth of the grid, forming pattern $p_{n}$, from the topmost row, which is the initial row in plane $z=1$. The first tile placed in each row follows a diagonal path starting from the second tile from the left. Growth of each row expands left and right from the tile along the diagonal. The bit value for each column is propagated south from the top column, and the bit value for each row is propagated from the diagonal tile that initiates that row. In this way, the $i$th row propagates the same bit value as the $i$th column.
Upon completion of the first row of $z=1$, the final module of the construction begins growth. This module consists of 52 tile types that grow south from that first row to make the full square in $z=1$ while copying the pattern downward and to both sides to form the grid pattern $p_{n}$. (See Figure 13 for an example.)
6.3 Correctness of proof
Throughout the definition of the construction, we have explained the correctness of each component, so in this section we summarize those arguments to complete the proof of Theorem 6.1. We first note that, by definition, our construction creates a pattern in the plane $z=1$ using 7 colors: 3 for the boundary rows and columns of each cell, and 4 for the interior regions of each cell. Which 3 boundary colors are used for any particular $p_{n}$ depends upon the result of the first system simulated. Next, we argue that for every $n$ the system simulates every possible aTAM system with $\leq n$ tile types and a single-tile seed, or a functionally equivalent aTAM system. The details of how the enumeration of all such systems is guaranteed are given in Section 6.1. While the enumeration will generate multiple equivalent systems (e.g. given two possible tile types $t_{1}$ and $t_{2}$, there will be two tile sets generated with exactly those two tile types, but listed in both orders, $\{t_{1},t_{2}\}$ and $\{t_{2},t_{1}\}$) that only serves to make the series of simulations, and bit sequence generated, longer. This results in $p_{n}$ differing from each of those multiply-counted systems in at least as many locations of each grid cell as the total number of times those systems are simulated. The fact that a bit is gathered for each simulation to guarantee that $p_{n}$ differs from it in at least one location is shown in Section 6.2. Thus, it is shown that $p_{n}$ must differ from the pattern made by all aTAM systems with $\leq n$ tile types. Finally, we just need to show that the tile complexity of the barely-3DaTAM system $\mathcal{T}_{p_{n}}$ is $O(\log{n}/\log{\log{n}})$. For this, we note that the tiles of all components are constant with respect to $n$, with the exception of the initial component that unpacks the value of $n$ from its optimal encoding using $O(\log{n}/\log{\log{n}})$ tile types, for an overall tile complexity of $O(\log{n}/\log{\log{n}})$.
Additional technical details, including pseudocode for the algorithms of the Turing machine $M$ and its simulations of all systems with $\leq n$ tile types, the layout of data structures used during the simulation of a system, and time complexity analysis, can be found in Section 7 of the technical appendix.
@@ 6.4 Extending a pattern $p_{n}$ to infinitely cover $\mathbb{Z}^{2}$
We now present a version of the previous result in which the patterns produced extend infinitely, rather than being bounded by squares.
Theorem 6.2
For all $n\in\mathbb{Z}^{+}$, there exists a 7-colored pattern, $p_{n_{\infty}}$, that infinitely covers the plane $\mathbb{Z}^{2}$ such that no aTAM system $\mathcal{T}_{\leq n}=(T,\sigma,\tau)$ where $|T|\leq n$, $|\sigma|=1$, and $\tau\leq 2$ weakly self-assembles $p_{n_{\infty}}$, but a barely-3DaTAM system $\mathcal{T}_{p_{n}}=(T_{p_{n}},\sigma_{p_{n}},2)$ where $|T_{p_{n}}|=O(\log{n}/\log{\log{n}})$ and $|\sigma_{p_{n}}|=1$ weakly self-assembles $p_{n_{\infty}}$.
To prove Theorem 6.2, we extend the construction from the proof of Theorem 6.1 so that every pattern $p_{n}$ from the proof of Theorem 6.1 is extended to infinitely cover the $\mathbb{Z}^{2}$ plane, becoming $p_{n_{\infty}}$, by usage of “grid-reconstruction,” i.e., a method of copying the square grid infinitely to each side. This requires $O(1)$ unique tile types, and its assembly may be used to replicate patterns of any length $n>1$, as infinitely tiling a single bit pattern is trivial. Due to the symmetry exhibited by all $m\times m$ squares along their northeast $\rightarrow$ southwest diagonal, copies of the same pattern may be copied along these diagonals infinitely.
Given that $p_{n}$ forms on the $z=1$ layer of an $m\times m$ square, once the initial square forms, this extension causes it to be branched off of in the four cardinal directions, starting from positions $(0,m-1)$ for southward and westward expansion, and $(m-1,0)$ for northward and eastward expansion. These patterns assemble new copies of themselves upon two of three boundaries defined along their edges with glue signals, which are either propagated along from the previously formed square, or are created upon the diagonal grid-generating signal making contact with an associated boundary.
Figure 14: (a) The growth of the grid, forming pattern $p_{north}$, identical to pattern $p_{n}$, with adjustments made for expansion to the north, from its bottom-most row, which is the initial row on the plane $z=1$. The first tile placed in each row follows a diagonal path, colored in red for clarity, starting from the first tile to the right. Growth of each row expands left and right from the tile along the diagonal. Upon the leftmost and rightmost edges are bounding tiles, colored in black for clarity, which denote the leftmost and rightmost bounds of $p_{north}$. Once the diagonal makes contact with the left bound, it generates a top-bound at height $h+1$, where $h$ is the height at which the diagonal makes contact with the left bound. Once the top-bound’s signal is propagated to reach the right-bound, a new grid-forming diagonal is created, starting the process over again. (b) The infinite growth of $m\times m$ squares illustrated at a high level, with the initial $m\times m$ square which generates $p_{n}$ colored in yellow, direction of growth colored in red, and direction of copy assembly, which utilizes the signals of those assembled above or below them colored in black.
This basic assembly is used in all four cardinal directions in order to form a cross of $m\times m$ squares, each of which extends infinitely. Once a column is formed in either the east or west direction, and a corresponding row is formed in the north or south direction, tiles containing the bit values of that column to their north or south, and rows to their east or west are able to propagate out along those patterns created in the cross assembly. This allows for the infinite tiling of the same $m\times m$ square infinitely across the $\mathbb{Z}^{2}$ plane to form $p_{n_{\infty}}$.
(a) Infinite tiling along the bit sequence $0110010010101$.
(b) Infinite tiling along the bit sequence $0110010010101$ with the cross assembly marked with red tiles along grid-creating diagonals, and black tiles along bounds.
Figure 15: Assemblies representing the infinite tiling of $m\times m$ squares corresponding to bit sequence $0110010010101$ across $\mathbb{Z}^{2}$. With creation and bounds marked, and unmarked.

### 10 Mar 2024 v2

@@ 6 Barely-3DaTAM patterns
For all $n\in\mathbb{Z}^{+}$, for some $m\in\mathbb{Z}^{+}$ there exists a 7-colored $m\times m$ pattern, $p_{n}$, such that (1) no aTAM system $\mathcal{T}_{\leq n}=(T,\sigma,\tau)$ exists where $|T|\leq n$, $|\sigma|=1$, and $\mathcal{T}_{\leq n}$ weakly self-assembles $p_{n}$, but (2) a barely-3DaTAM system $\mathcal{T}_{p_{n}}=(T_{p_{n}},\sigma_{p_{n}},2)$ exists where $|T_{p_{n}}|=O(\log{n}/\log{\log{n}})$, $|\sigma_{p_{n}}|=1$, and $\mathcal{T}_{p_{n}}$ weakly self-assembles $p_{n}$.
@@
Figure 6: Example $p_{n}$ patterns created by the construction in the proof of Theorem 6.1. The (repeatedly copied) binary sequence derived from the results of the simulations of aTAM systems starts at the top, with the two colors of that row, and all subsequent boundary rows, being determined by the first bit of that pattern. During the downward growth from that row, during which the full $m\times m$ square is formed, the repeating grid formed by the copies of that pattern is copied both downward and to both sides. The boundary columns also have two colors determined by the first bit of the sequence (one of them the same as in the boundary rows) for a total of 3 boundary colors. The interiors always use the same 4 colors.
Proof sketch
@@
The bit sequence that determines the colors used by the boundaries, and the ordering of the colors on the boundaries and in the interiors, is determined via simulations of a series of aTAM systems. Intuitively, our proof utilizes a construction that performs a diagonalization against all possible aTAM systems with $\leq n$ tile types by simulating each for a bounded number of steps, and for each keeping track of the color of tile it places in a location specific to the index of that system so that it can ultimately generate the colored pattern $p_{n}$ that differs in at least one location from every simulated system.
Figure 7: Schematic overview of the portion of the construction for the proof of Theorem 6.1 that grows in plane $z=0$. Modules are not shown to scale. (Green) The seed tile, (Fuchsia) the base conversion module that unpacks the binary representation of $n$, (Aqua) the module that simulates the Turing machine $M$ on input $n$, which itself simulates each aTAM system with $\leq n$ tile types in sequence and saves a result bit (Black or White) for each, (Yellow and Grey) the pattern of result bits is copied repetitively to the right until it covers the entire top row. (Light Grey) A “filler” tile causes the assembly to form a complete square.
The dimensions of each $c\times c$ cell are $c=\texttt{SF}(n)$, where $\texttt{SF}(n)$ is a function that takes a number of tile types and returns an upper bound on the number of all possible singly-seeded aTAM systems with $\leq n$ tile types and $\leq 8$ colors. (Note that $\texttt{SF}(n)$ is actually greater than the number of such systems, and details of $\texttt{SF}(n)$ can be found in Section 10.) The colors of the rows and columns encode the bit sequence generated by the simulations, with the same bit sequence encoded in both the rows and the columns via an assignment of colors. There is a unique color assigned for each intersection of two bits (i.e. 00, 01, 10, and 11), with 4 colors reserved for boundaries of grid cells and 4 separate colors reserved for the interior locations of the grid cells. Therefore, the colored pattern of every cell differs from the assemblies produced by all aTAM systems with $\leq n$ tile types. It forms on the top layer of a two-layered $m\times m$ square where $m=O(n^{21n})$, and the barely-3DaTAM system that forms it uses $O(\log{n}/\log{\log{n}})$ tiles since the tiles types for all modules are constant except those that encode $n$ using optimal encoding [1]. (Note that the value of $m$ could be smaller, $O(n^{4n}n^{8})$, if it wasn’t desired that both planes be the same size. Additionally, by having $M$ simulate systems with larger tile sets, the value of $m$ would increase but the difference in tile complexity between the barely-3DaTAM system and the systems incapable of making its patterns could be increased beyond the current exponential bound.) Additional technical details, including pseudocode for the algorithms of the Turing machine $M$ and its simulations of all systems with $\leq n$ tile types, the layout of data structures used during the simulation of a system, and time complexity analysis, can be found in Section 10 of the technical appendix.
@@ 6.1 Extending a pattern $p_{n}$ to infinitely cover $\mathbb{Z}^{2}$
Although it is already known that there are infinite patterns that can’t weakly self-assemble from any finite-sized tile set [12], the following corollary simply shows how the previously defined patterns can be extended to infinitely cover the plane, while keeping an arbitrary spread in the tile complexity required by aTAM and barely-3DaTAM systems.
Corollary 1
For all $n\in\mathbb{Z}^{+}$, there exists a 7-colored pattern, $p_{n_{\infty}}$, that infinitely covers the plane $\mathbb{Z}^{2}$ such that (1) no aTAM system $\mathcal{T}_{\leq n}=(T,\sigma,\tau)$ exists where $|T|\leq n$, $|\sigma|=1$, and $\mathcal{T}_{\leq n}$ weakly self-assembles $p_{n_{\infty}}$, but (2) a barely-3DaTAM system $\mathcal{T}_{p_{n}}=(T_{p_{n}},\sigma_{p_{n}},2)$ exists where $|T_{p_{n}}|=O(\log{n}/\log{\log{n}})$, $|\sigma_{p_{n}}|=1$, and $\mathcal{T}_{p_{n}}$ weakly self-assembles $p_{n_{\infty}}$.
To prove Corollary 1, we extend the construction from the proof of Theorem 6.1 so that every pattern $p_{n}$ from the proof of Theorem 6.1 is extended to infinitely cover the $\mathbb{Z}^{2}$ plane, becoming $p_{n_{\infty}}$, by usage of “grid-reconstruction,” i.e., a method of copying the square grid infinitely to each side. This requires $O(1)$ unique tile types in addition to those used in the previous construction. Due to the symmetry exhibited by all $m\times m$ squares of $p_{n}$ patterns along their northeast $\rightarrow$ southwest diagonal, copies of the same pattern may be copied along these diagonals infinitely. Details of the construction can be found in Section 11.